# Position: On-Premises LLM Deployment Demands a Middle Path: Preserving Privacy Without Sacrificing Model Confidentiality

### v2

@@ 5 Experiments
In this section, we conduct experiments to validate the correctness of our positions. Specifically, we answer the following research questions:
  • RQ1. Can query-based distillation attacks distill the functionality of the entire model under the baseline deployment that secures the top layer?
@@
  • RQ3. Does securing bottom layers (SOLID) offer a better balance between model theft risk and customization performance compared to baseline deployments?
@@ 5.1 Experimental Settings
Attack Methods. We distill models produced by different protection approaches using three attack methods: FT-all, FT-closed and SEM. Following (He et al., 2021), a diverse attack set is required for full distillation. Therefore, we merge data evenly from two general datasets, MMLU (Hendrycks et al., 2021) and Alpaca 52k (Wang et al., 2022), resulting in a 51k combined set.
Baselines. We evaluate the effectiveness of SOLID by comparing it with three baselines: DarkneTZ (Mo et al., 2020), SAP-DP, and the fully-secured approach (Eiras et al., 2024). DarkneTZ protects only the final decoder layer while the SAP (Shen et al., 2023) framework exposes the first six decoder layers and secures the rest. SAP-DP extends SAP by adding Laplace noise to model outputs to enhance protection (Lee et al., 2018). The fully-secured approach represents the extreme, securing all layers for maximal security.
Implementation Details of SOLID. We apply the SOLID algorithm to identify the smallest secure set $I$ such that $R(I)\leq(1+\varepsilon)R([L])$. To calculate distillation difficulty (DD), we use cross-entropy loss and approximate the expectation over samples distributed on the general domain and randomly initialized secured parameters. This is done using a 1,500-sample evaluation set randomly sampled from the MMLU and Alpaca 52k, with secured parameters initialized via Xavier initialization and averaged over three random seeds (20, 42, 1234). In our experiments, we find that $\varepsilon=0.05$ yields optimal performance.
Table 1: Distillation ratios on 6 functionalities under FT-all (SOLID||||SAP-DP||||Fully-secured||||DarkneTZ). “H.E.” in Code domain presents the benchmark “HumanEval”. More results and details are available in Appendix C.2.
@@
Rsn. PIQA 62.6||||59.8||||63.0||||99.3 64.7||||64.7||||64.6||||99.1 63.0||||61.2||||60.2||||92.2 68.3||||65.6||||65.7||||99.1
Winogrande 68.5||||67.7||||68.3||||98.3 76.8||||74.8||||76.6||||100. 67.2||||69.0||||68.3||||89.5 68.3||||64.9||||64.8||||99.1
ARC-easy 31.9||||32.8||||31.3||||98.5 36.3||||35.5||||34.9||||97.6 32.3||||34.7||||32.0||||86.6 43.2||||35.3||||33.9||||99.5
ARC-challenge 38.5||||38.1||||44.2||||99.2 47.8||||46.6||||50.9||||100. 39.7||||42.6||||44.5||||81.4 36.8||||36.6||||35.3||||99.5
Hellaswag 31.4||||31.4||||32.4||||98.1 33.9||||34.0||||35.0||||96.6 32.2||||32.0||||31.3||||84.6 37.4||||37.3||||34.3||||96.5
Read. LAMBADA 0.01||||0.00||||0.00||||88.6 0.02||||0.00||||0.01||||92.2 0.16||||0.00||||0.01||||67.9 1.34||||0.04||||0.00||||94.6
BoolQ 47.2||||47.1||||53.9||||100. 59.5||||56.0||||65.0||||99.6 48.3||||46.8||||56.7||||97.3 56.7||||50.3||||55.8||||100.
SQuADv2 1.50||||1.68||||0.34||||55.3 0.68||||0.88||||0.82||||59.5 1.69||||0.36||||0.93||||50.7 3.65||||0.39||||0.90||||62.9
OBQA 54.5||||54.5||||57.1||||99.6 57.4||||52.5||||59.2||||94.8 57.7||||56.8||||56.3||||84.0 0.00||||0.00||||0.02||||94.3
Knl. NaturalQuestions 0.00||||0.02||||0.00||||40.1 0.01||||0.01||||0.08||||53.6 0.00||||0.00||||0.02||||31.8 0.01||||0.00||||0.06||||87.4
TriviaQA 0.00||||0.02||||0.00||||72.3 0.00||||0.00||||0.03||||73.8 0.00||||0.00||||0.01||||38.7 0.01||||0.00||||0.01||||68.9
Code MBPP&H.E. 0.00||||0.00||||0.00||||58.6 0.00||||0.00||||0.00||||90.9 0.00||||0.00||||0.00||||40.2 0.00||||0.00||||0.00||||91.1
Math GSM8K 0.02||||0.00||||0.06||||79.6 0.00||||0.00||||0.00||||78.6 0.00||||0.00||||0.00||||31.1 0.00||||0.00||||0.00||||86.2
Gen. MMLU 36.8||||38.3||||36.5||||96.7 52.9||||50.0||||53.3||||110. 40.4||||36.9||||37.2||||81.7 42.6||||40.3||||40.5||||99.5
BBH 0.00||||0.00||||0.00||||93.3 0.00||||0.00||||0.00||||101. 0.00||||0.00||||0.00||||63.3 0.01||||0.00||||0.00||||94.8
Average Distillation Ratio($\downarrow$) 21.9||||21.8||||22.8||||77.9 25.3||||24.4||||25.9||||86.5 22.5||||22.4||||22.8||||73.7 23.9||||22.3||||22.4||||88.9
Secured Ratio($\downarrow$) 2.50||||92.5||||100.||||1.25 3.16||||81.3||||100.||||3.16 3.16||||81.3||||100.||||3.16 6.25||||81.3||||100.||||3.16
Evaluation Benchmarks We assess adaptability on six downstream tasks: Code (Zheng et al., 2024b), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023b), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). To fully evaluate recovered functionalities, we focus on six capabilities domains following Llama2 report (Touvron et al., 2023). Specifically, we assess the recovered model across sixteen benchmarks grouped into (1) Commonsense Reasoning (Rsn.); (2) Reading Comprehension (Read.); (3) World Knowledge (Knl.); (4) Code; (5) Math; and (6) General Ability (Gen.).
Metrics. We measure customization through model’s performance on benchmarks. For security, we calculate the “Average Distillation Ratio” (ADR) by averaging the distillation ratios across benchmarks. A lower ADR indicates higher security offered by the secure set.
@@ 5.2 Failure in Defense (RQ1)
We evaluate the security of DarkneTZ using three distillation strategies. Based on the results shown in Tables 1 and 2, we have following observations.
Observation 1: DarkneTZ, which secures only the last decoder layer, fails to protect the model against all three attacks (Table 1). DarkneTZ achieves ADRs generally exceeding 73% across four models, significantly higher than the average 24% ADR obtained by the fully-secured approach that offers maximum security. Specifically, on Llama2-70B and Phi-2, the distillation ratio for the Reasoning (Rsn.) task consistently exceeds 98%, indicating near-complete replication of model functionality. On Llama2-7B, the attack achieves a distillation ratio over 100% on the MMLU and BBH, suggesting the distilled model even outperforms the victim on these tasks. Similarly, Table 2 highlights consistent failure patterns against FT-closed and SEM attacks, with ADRs consistently above 75%. This confirms that securing only the top layers is insufficient to prevent the replication of model functionality.
@@
5.3 Security-Customization Trade-off (RQ2)
We conduct two experiments to analyze the impact of secured layer placement and quantity on the trade-off between security and customization. First, we secure one layer in Llama2-7B and two in Phi-2, varying their placement. Second, we incrementally secure both models by adding protected layers, starting from the smallest module (k_project) of the first decoder layer. These models are evaluated under the FT-all distillation attack and customized for the math domain. The results, as shown in Figure 4, lead to the following observations. More results and details can be found in Appendix B.7 and B.8.
Observation 2: Secured layer placement significantly impacts security, consistent with Theorem 1, but has limited effect on customization performance (Figure 4(a)). For Llama2-7B, security transitions at the fourteenth layer, with ADR consistently near 20% for earlier sets, indicating stronger security than protecting later layers. Meanwhile, customization accuracy remains stable across placements, highlighting the advantage of securing pre-transition layers. In contrast, Phi-2 transitions earlier at the first layer set, where only the first set balances security and customization, with later sets reducing security. These results suggest that securing layers before the transition layer optimizes the security-customization trade-off.
Observation 3: Increasing the number of secured layers enhances security but reduces customization (Figure 4(b)). The ADR of Llama2-7B decreases from 85% to 22% after securing an entire decoder layer, indicating improved security. However, customization accuracy drops from 29% to 21% as the number of secured layers increases from one to five, reflecting reduced customization flexibility. A similar trend is observed in Phi-2, suggesting that while increasing the number of secured layers enhances security (lower ADR), it negatively impacts customization flexibility (lower ACC) in both models.
5.4 Effectiveness of Securing Bottom Layers (RQ3)
Table 2: Distillation ratios of Llama2-70B under FT-closed (FT-c.) and SEM attacks.
@@ 5.4 Effectiveness of Securing Bottom Layers (RQ3)
Observation 4: SOLID achieves security against model distillation comparable to fully secured protection while securing significantly fewer parameters (Table 1). SOLID attains a similar security level (ADR) to SAP-DP and the fully secured approach across four architectures and multiple domains, securing at most 6.25% of parameters—compared to at least 80% for the others. For instance, on Llama2-70B, SOLID secures only 1.25% of parameters yet achieves an ADR of 21.9%, comparable to SAP-DP (21.8%) and the fully secured approach (22.8%). Additionally, SOLID maintains this security under FT-closed and SEM attacks, with ADR differences from SAP-DP and the fully secured approach remaining below 2.1% across six domains (Table 2). Details are in Appendix C.4 and C.5.
@@
Observation 5: SOLID approaches the customization flexibility of full parameter fine-tuning, consistently outperforming baseline deployments (Figure 5). We observe that the performance of SOLID consistently matches full parameter fine-tuning across four architectures, with differences within 4%. Additionally, in the Law domain, SOLID improves performance by 10% over SAP-DP and the fully-secured approach on Llama2-70B, with a 35% improvement on the 7B models. Similar gains are evident on Phi-2, though the improvement in the Law domain reduces to 1%. These results demonstrate that securing a small portion of parameters minimally impacts customization, while providing strong protection against distillation attacks. Further results can be found in Appendix B.6 and C.3.
6 Alternative Views
Acknowledging diverse views and potential curiosities regarding LLM deployment design, we objectively examine several alternate viewpoints.
  • Alternative: Enhanced Regulation. Enhanced government data regulations can partially mitigate ethical privacy concerns and strengthen user confidence in data uploads and fine-tuning APIs, thereby promoting industry development. However, they do not eliminate the risk of data breaches and may struggle to address adaptive threats, rendering them insufficient for organizations with stringent privacy requirements, such as government agencies.
  • Alternative: Hybrid Privacy-Aware Deployment. Federated hybrid deployment enables users to deploy early model layers and process sensitive data locally while hosting later layers in the cloud for enhanced protection (Shen et al., 2023). It employs privacy-preserving techniques (e.g., differential privacy) to sanitize inputs before generating intermediate representations, ensuring secure data transmission. This approach preserves customization performance while improving both data privacy and model security. However, supervised fine-tuning still requires data uploads, and intermediate exchanges remain vulnerable to breaches (Zhou et al., 2022), thereby compromising the overall security assurance.
  • Alternative: Advanced Open-Weight Models. The success of DeepSeek-R1 (Guo et al., 2025) has showcased the power of open-weight models, raising expectations for more robust open-source alternatives. Developing open-weight models with strong security alignment could offer a viable solution to the ongoing dilemma between data privacy and model confidentiality. These models, capable of ensuring both usability and robust misuse detection, could address ethical concerns and serve as an alternative to the privacy-preserving deployment methods. However, their adoption may be constrained by limited profitability.

### v3

@@ 4 Experiments
In this section, we conduct experiments to answer the following research questions:
  • RQ1. Can query-based distillation attack distill the functionality of the entire model under the baseline deployment that secures the top layer?
@@
  • RQ3. Does SOLID offer a better balance between model theft risk and customization performance compared to baseline deployments?
  • RQ4. How does SOLID optimize this trade-off? Is it effective for both large and small models?
@@ 4.1 Experimental Settings
Attack Methods. We distill models produced by different protection approaches using three attack methods: FT-all, FT-closed and SEM. Following (He et al., 2021), a diverse attack set is required for full distillation. Therefore, we merge data evenly from two general datasets, MMLU benchmark (Hendrycks et al., 2021) and Alpaca 52k (Wang et al., 2022), resulting in a 51k combined set. Additionally, we build four larger general datasets (100k–500k) to strengthen the attack.
Baselines. We compare SOLID with three baselines: SAP-DP, the fully-secured approach (Eiras et al., 2024), and DarkneTZ (Mo et al., 2020). The SAP (Shen et al., 2023) framework exposes the first six decoder layers and secures the rest. SAP-DP extends SAP by adding Laplace noise to model outputs to enhance protection (Lee et al., 2018). The fully-secured approach represents the extreme, securing all layers for maximal security, while DarkneTZ protects only the final decoder layer.
Implementation Details of SOLID. We apply the SOLID algorithm to identify the smallest secure set $I$ such that $R(I)\leq(1+\varepsilon)R([L])$. To calculate distillation difficulty (DD), we use cross-entropy loss and approximate the expectation over samples distributed on the general domain and randomly initialized secured parameters. This is done using a 1,500-sample evaluation set randomly sampled from the MMLU benchmark and Alpaca 52k, with secured parameters initialized via Xavier initialization and averaged over three random seeds (20, 42, 1234). In our experiments, we find that $\varepsilon=0.05$ yields optimal performance.
Evaluation Benchmarks We assess adaptability on six downstream tasks: Code (Zheng et al., 2024a), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023a), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). To fully evaluate recovered functionalities, we focus on six capabilities domains following Llama2 report (Touvron et al., 2023). Specifically, we assess the recovered model across sixteen benchmarks grouped into (1) Commonsense Reasoning (Rsn.); (2) Reading Comprehension (Read.); (3) World Knowledge (Knl.); (4) Code; (5) Math; and (6) General Ability (Gen.).
Metrics. We measure customization through model’s improvements on benchmarks. For security, we calculate the “Average Distillation Ratio” (ADR) by averaging the distillation ratios across benchmarks. A lower ADR indicates higher security offered by the secure set.
@@
Rsn. PIQA 62.6||59.8||63.0||99.3 64.7||64.7||64.6||99.1 63.0||61.2||60.2||92.2 68.3||65.6||65.7||99.1
Winogrande 68.5||67.7||68.3||98.3 76.8||74.8||76.6||100. 67.2||69.0||68.3||89.5 68.3||64.9||64.8||99.1
ARC-easy 31.9||32.8||31.3||98.5 36.3||35.5||34.9||97.6 32.3||34.7||32.0||86.6 43.2||35.3||33.9||99.5
ARC-challenge 38.5||38.1||44.2||99.2 47.8||46.6||50.9||100. 39.7||42.6||44.5||81.4 36.8||36.6||35.3||99.5
Hellaswag 31.4||31.4||32.4||98.1 33.9||34.0||35.0||96.6 32.2||32.0||31.3||84.6 37.4||37.3||34.3||96.5
Read. LAMBADA 0.01||0.00||0.00||88.6 0.02||0.00||0.01||92.2 0.16||0.00||0.01||67.9 1.34||0.04||0.00||94.6
BoolQ 47.2||47.1||53.9||100. 59.5||56.0||65.0||99.6 48.3||46.8||56.7||97.3 56.7||50.3||55.8||100.
SQuADv2 1.50||1.68||0.34||55.3 0.68||0.88||0.82||59.5 1.69||0.36||0.93||50.7 3.65||0.39||0.90||62.9
OBQA 54.5||54.5||57.1||99.6 57.4||52.5||59.2||94.8 57.7||56.8||56.3||84.0 0.00||0.00||0.02||94.3
Knl. NaturalQuestions 0.00||0.02||0.00||40.1 0.01||0.01||0.08||53.6 0.00||0.00||0.02||31.8 0.01||0.00||0.06||87.4
TriviaQA 0.00||0.02||0.00||72.3 0.00||0.00||0.03||73.8 0.00||0.00||0.01||38.7 0.01||0.00||0.01||68.9
Code MBPP&H.E. 0.00||0.00||0.00||58.6 0.00||0.00||0.00||90.9 0.00||0.00||0.00||40.2 0.00||0.00||0.00||91.1
Math GSM8K 0.02||0.00||0.06||79.6 0.00||0.00||0.00||78.6 0.00||0.00||0.00||31.1 0.00||0.00||0.00||86.2
Gen. MMLU 36.8||38.3||36.5||96.7 52.9||50.0||53.3||110. 40.4||36.9||37.2||81.7 42.6||40.3||40.5||99.5
BBH 0.00||0.00||0.00||93.3 0.00||0.00||0.00||101. 0.00||0.00||0.00||63.3 0.01||0.00||0.00||94.8
Average Distillation Ratio($\downarrow$) 21.9||21.8||22.8||77.9 25.3||24.4||25.9||86.5 22.5||22.4||22.8||73.7 23.9||22.3||22.4||88.9
Secured Ratio($\downarrow$) 2.50||92.5||100.||1.25 3.16||81.3||100.||3.16 3.16||81.3||100.||3.16 6.25||81.3||100.||3.16
Table 1: Distillation ratios across six functionalities under FT-all (SOLID||SAP-DP||Fully-secured||DarkneTZ). “H.E.” in the Code domain denotes the benchmark HumanEval. Green and red indicate the overall best- and worst-performing methods, respectively. Additional results are provided in Appendix C.2.
@@ 4.2 Failure in Defense (RQ1)
We evaluate security of DarkneTZ using three distillation strategies. Based on the results shown in Tables 1 and 2, we have following observations.
Obs1: DarkneTZ, which secures only the last decoder layer, fails to protect the model against all three attacks. As shown in Table 1, DarkneTZ achieves ADRs generally exceeding 73%. Notably, on Llama2-7B, it surpasses 100% distillation ratio on the MMLU and BBH datasets, indicating that the distilled model outperforms the original on these tasks. Similarly, Table 2 highlights consistent failure patterns against FT-closed and SEM attacks, with DarkneTZ maintaining ADRs above 75%, demonstrating the ability of these strategies to recover significant model functionality.
4.3 Security-Customization Trade-off (RQ2)
@@
We conduct two experiments to analyze the impact of secured layer placement and quantity on the trade-off between security and customization. First, we secure one layer in Llama2-7B and two in Phi-2, varying their placement. Second, we incrementally secure both models by adding protected layers, starting from the smallest module (k_project) of the first decoder layer. These models are evaluated under the FT-all distillation attack and customized for the math domain. The results, as shown in Figure 4, lead to the following observations.
Obs2: Secured layer placement significantly impacts security, consistent with Theorem 1, but has small effect on customization performance. As shown in Figure 4(a), for Llama2-7B, security transitions at the fourteenth layer, with ADR consistently near 20% for earlier sets, indicating stronger security than protecting later layers. Meanwhile, customization accuracy remains stable across placements, highlighting the advantage of securing pre-transition layers. In contrast, Phi-2 transitions earlier at the first layer set, where only the first set balances security and customization, with later sets reducing security. These results suggest that securing layers before the transition layer optimizes the security-customization trade-off. Results for Mistral-7B and Phi-1.5 are in Appendix B.7.
Obs3: Increasing the number of secured layers enhances security but reduces customization. As shown in Figure 4(b), the ADR of Llama2-7B decreases from 85% to 22% after securing an entire decoder layer, indicating improved security. However, customization accuracy drops from 29% to 21% as the number of secured layers increases from one to five, reflecting reduced customization flexibility. A similar trend is observed in Phi-2, suggesting that while increasing the number of secured layers enhances security (lower ADR), it negatively impacts customization flexibility (lower ACC) in both models. Further details are in Appendix B.8.
@@
Table 2: Distillation ratios of Llama2-70B under FT-closed and SEM attacks.
4.4 Effectiveness of SOLID (RQ3)
Obs4: SOLID offers comparable security against model distillation to the highest level of protection (fully-secured), while securing significantly fewer parameters. As shown in Table 1, SOLID achieves a similar security level (ADR) to SAP-DP and the fully-secured approach across four architectures and various domains, while securing at most 6.25% of parameters, compared to at least 80% for the others. For example, on Llama2-70B, SOLID secures only 1.25% of parameters yet achieves an ADR of 21.9%, comparable to SAP-DP (21.8%) and the fully-secured approach (22.8%), which protect 92.5% and 100% of parameters, respectively. Furthermore, under FT-closed and SEM attacks, SOLID also matches the security level provided by SAP-DP and the fully-secured approach. Table 2 shows that under FT-closed attack, the ADR differences between SOLID, SAP-DP, and the fully-secured approach remain below 2.1% across six domains. Similarly, under SEM attack, the distillation ratios closely aligned with the other two approaches. These results confirm that SOLID effectively protects against distillation attacks while securing significantly fewer parameters. More details are in Appendix C.4 and C.5.
Scale Rsn. Read. Knl. C.&M. Gen. ADR ADR-Da.
51k 51.7 21.6 0.01 0.00 28.3 25.3 86.5
100k 51.3 21.5 0.13 0.00 29.6 25.3 89.1
200k 51.4 21.7 0.11 0.00 29.7 25.2 91.3
300k 51.6 21.7 0.11 0.00 30.5 25.5 94.5
500k 51.8 22.0 0.09 0.00 30.8 25.8 96.9
Table 3: SOLID vs. Dataset scales. ADR-Da. represents the ADR by DarkneTZ. Details are in Appendix C.6.
Obs5: The security of SOLID cannot be easily compromised by simply increasing the dataset scale. As shown in Table 3, the distillation ratios for SOLID increase marginally with larger datasets, showing only a 0.5% ADR rise when scaling from 51k to 500k samples. In contrast, DarkneTZ exhibits a significant increase in the ADR, from $86.5\%$ to $96.9\%$, over the same dataset size range. This highlights the robustness of SOLID’s security against increasing attack dataset sizes. Details of the attack datasets are provided in Appendix 5.
@@ 4.4 Effectiveness of SOLID (RQ3)
Figure 6: ADRs vs. average customization score. Points closer to the bottom-right indicate better balance.
@@
Obs6: SOLID consistently outperforms baseline deployments in customization while achieving security levels comparable to fully-secured approaches. Its customization performance closely matches full parameter fine-tuning. As shown in Figure 5, SOLID improves scores in the Law domain by 10% over SAP-DP and fully-secured methods on Llama2-70B, and by 35% on 7B models. Similar trends are observed on Phi-2, though the gain in Law narrows to 1%. Additionally, the performance of SOLID consistently matches the performance of full parameter fine-tuning across four architectures, with differences within 4%. This indicates that securing a small subset of parameters preserves customization while ensuring strong protection against distillation attacks. Further results are in Appendix B.6 and C.3.
We summarize the security and customization performance of each deployment in Figure 6. SOLID achieves an optimal balance between distillation prevention and customization, outperforming other baselines. In the next subsection, we discuss how the distillation difficulty metric optimizes the security-customization trade-off.
4.5 Discussion on DD (RQ4)
We assess the efficacy of distillation difficulty (DD) in estimating distilled model performance by calculating the Pearson and Spearman correlation coefficients between DD and ADR across different domains. To address uncertainties regarding the effectiveness of DD in shallow transformers, we secure and attack 2-layer secured sets in OPT-350M (Zhang et al., 2022), which has 350M parameters. Based on the results shown in Figure 7, we have following observations.
Obs7: DD is effective in larger models, with a clear negative correlation between DD and average distillation ratios. As shown in Figure 7 (a), the Pearson coefficient for Llama2-7B consistently remains below -0.80, reaching as low as -0.98. We also observe similar phenomena in other models with varying architectures and sizes, confirming DD as a reliable predictor of distilles model performance and the effectiveness of SOLID. Results of Spearman coefficients are in Appendix B.9.
Obs8: DD is ineffective in smaller OPT model, with notable inconsistencies with ADRs. As shown in Figure 7 (b), DD exhibits weak negative correlation with ADR in OPT-350M (coefficients > -0.33), showing its unsuitability for predicting distillation performance. Additionally, optimal security is achieved by protecting the middle layers rather than the initial or output layers, making SOLID unable to identify the smallest secured set. Further details are provided in Appendix C.9.
5 Related Works
On-premises deployment. Using LLM services for customization poses significant privacy risks, as user data may be exposed during transmission, storage, and processing Li et al. (2024c). To mitigate this, privacy-sensitive sectors require on-premises deployment of LLMs, which retains both data and models within their local infrastructure Schillaci (2024); Nevo et al. (2024). However, this shifts security risks to vendors, who lose control over model use and face increased threats of theft, especially from hardware- and communication-based attacks on GPUs Nayan et al. (2024); Rakin et al. (2022). To secure models locally, hardware-based protections such as TrustZone have been proposed Pinto and Santos (2019); Zhang et al. (2024a); Li et al. (2024a), but they are resource-intensive with limited flexibility Mo et al. (2020). A more adaptable approach is layer-wise security, which protects only selected layers Lin et al. (2024); Chen et al. (2024); Zhang et al. (2024b). While prior work suggests securing shallow Elgamal and Nahrstedt (2020), intermediate Shen et al. (2022), or output layers Huang et al. (2024), most studies focus on smaller models. Our results show that securing a few, well-chosen bottom layers of LLMs can enhance security while preserving fine-tuning flexibility for on-premises deployment.
Model Distillation Attacks. Model distillation attacks allow adversaries to replicate model functionality using only black-box access, a process also known as functional extraction (Nevo et al., 2024; Xu et al., 2024a; Ezzeddine et al., 2024). While distillation attacks have been extensively studied in smaller models, such as CNNs Orekondy et al. (2018), BERT Sanh et al. (2020); Zanella-Beguelin et al. (2021), and ReLU-based models Canales-Martínez et al. (2024); Jagielski et al. (2020), their effectiveness against LLMs remains an open question. Our work extends these attacks to Llama2-70B and demonstrates that securing only the output layer remains insufficient to prevent near-complete functionality replication.